OpenStudy (anonymous):

how does one simplify (3+sqrt(3))/(3-sqrt(3)) to 2+sqrt(3)?

OpenStudy (mathstudent55):

To simplify this fraction, \( \dfrac{3+\sqrt{3} }{3- \sqrt{3}} \) multiply it by \( \dfrac{3+\sqrt{3} }{3+ \sqrt{3}} \)

OpenStudy (mathstudent55):

In general, if you have a fraction with denominator \(a + \sqrt{b} \), and you need to rationalize the fraction, multiply the fraction by \( \dfrac{a - \sqrt{b} }{a - \sqrt{b}} \) This works because \( (a + \sqrt{b})(a - \sqrt{b}) \) \( = a^2 - (\sqrt{b})^2 \) \(= a^2 - b\) If a and b are rational, then \( a^2 - b\) is also rational.

OpenStudy (anonymous):

Got it, thanks!

OpenStudy (mathstudent55):

wlcm